Find the absolute extrema of f(x, y) = x^(2).y − 2.x + y on the
closed triangle T with vertices (1, 0),(5, 0) and (1, 4).
1
Expert's answer
2020-04-08T16:50:29-0400

"\frac {\partial f(x,y)} {\partial x}=2xy-2"

"\frac {\partial f(x,y)} {\partial y}=x^2+1"

We should solve the system "\begin {cases}
\frac {\partial f(x,y)} {\partial x}=0 \\
\frac {\partial f(x,y)} {\partial y}=0
\end {cases}" "\iff" "\begin {cases}
2xy -2=0 \\
x^2+1=0
\end {cases}"

This system has no solution. It means that there is no extremum point at all.


Our triangle T is bounded by these lines: "(x=1; y=-x+5; y=0)"

We must check how the surface behaves on these lines (triangle borders).

  1. If "y=0" then "f=-2x" is a straight line. It has no extremum point.
  2. If "x=1" then "f=2y-2" is a straight line too.
  3. If "y=-x+5" then "f=-x^3+5x^2-3x+5"

"f'=-3x^2+10x-3"

"f'=0" when "x=1\/3" or "x=3"

"x=1\/3" does not lie in the region bounded by the triangle.

if "x=3" then "y=2"

Now we have the following picture:


M1, M2, M3 are triangles vertices and the point M4 with coordinates (3,2) .

Let's find values of "f(x,y)" at these points.

M1: "f(1, 0) = -2"

M2: "f(5, 0) = -10"

M3: "f(1, 4)=6"

M4: "f(3, 2)= 14"


Finally, "\max \underset {T}{f(x,y)}=f(3,2)=14," "\min \underset {T}{f(x,y)}=f(5,0)=-10"